### v2

# A set $\mathcal{S}\subset\mathbb{N}$ is said to be a subset-sum-distinct or dissociated if all of its finite subsets have different sums. Alternately, an equivalent classification is if any equality of the form

Abstract
A set $\mathcal{S}\subset\mathbb{N}$ is said to be a subset-sum-distinct or dissociated if all of its finite subsets have different sums. Alternately, an equivalent classification is if any equality of the form
$\sum_{s\in\mathcal{S}}\varepsilon_{s}\cdot s=0$
where $\varepsilon_{s}\in\{-1,0,+1\}$ implies that all the $\varepsilon_{s}$’s are 0. For a dissociated set $\mathcal{S}$, we prove that for $c_{\ast}=\frac{1}{2}\log_{2}\left(\frac{\pi}{2}\right)$ and any $c_{\ast}-1<C<c_{\ast}$, we have
$\mathcal{S}(n)\,:=\,\mathcal{S}\cap[1,n]\,\leq\,\log_{2}n+\frac{1}{2}\log_{2}\log_{2}n+C$
for all $n\in\mathcal{N}_{C}$ with asymptotic density $\mathbf{d}\left(\mathcal{N}_{C}\right)=2-2^{c_{\ast}-C}$. Further, we consider the greedy algorithm for generating these sets and prove that this algorithm always eventually doubles. Finally, we also consider some generalizations of dissociated sets and prove similar results about them.
The Greedy Algorithm for Dissociated Sets
Sayan Dutta
sayandutta345@gmail.com
https://sites.google.com/view/sayan-dutta-homepage
1 Introduction
An $m$-element set $\mathcal{S}\subset\mathbb{N}$ is said to have distinct subset sums if all the $2^{m}$ subsets of $\mathcal{S}$ have different sums [11, 13, 12, 15, 14, 17, 30, 5, 8, 23, 24]. These sets have often been called dissociated sets in literature where they have been studied in more general settings [29, 6, 7, 21, 27, 25, 28]. Although this note is only concerned with subsets of $\mathbb{N}$, this is the terminology that will be used. An infinite subset of $\mathbb{N}$ is called dissociated if all its finite subsets are. Another way to describe these sets is
Definition 1.1.
A set $\mathcal{S}\subset\mathbb{N}$ is called dissociated if any equality of the form
$\sum_{s\in\mathcal{S}}\varepsilon_{s}\cdot s=0$
where $\varepsilon_{s}\in\{-1,0,+1\}$ implies that all the $\varepsilon_{s}$’s are 0.
An example of such sets is the set of powers of $2$.
Now, for such a set $\mathcal{S}$, let us define $\mathcal{S}(n):=\mathcal{S}\cap[1,n]$ to be the elements of $\mathcal{S}$ that are less than or equal to $n$. The first observation that was made in this field is the following.
Theorem 1.2 (Erdős).
For a dissociated set $\mathcal{S}$, we have $\left|\mathcal{S}(n)\right|\leq\log_{2}n+\mathcal{O}(\log_{2}\log_{2}n)$.
Proof.
Consider the map
$f:\mathcal{P}\left(\mathcal{S}(n)\right)\to\mathbb{N}$
that takes a subset of $\mathcal{S}(n)$ to the sum of its elements. Since $\mathcal{S}$ is dissociated, this map is injective. So, $2^{\left|\mathcal{S}(n)\right|}\leq n\cdot\left|\mathcal{S}(n)\right|+1$, hence completing the proof.
Theorem 1.2 was later improved to
$\left|\mathcal{S}(n)\right|\,<\,\log_{2}n\,+\,\frac{1}{2}\log_{2}\log_{2}n\,-\,\log_{2}c^{\ast}\,+\,o(1)$
where the constant $c^{\ast}$ has been improved subsequently to $1/4$ (Erdős and Moser [11]), $2\cdot(27)^{-1/2}$ (Alon and Spencer [2]), $\pi^{-1/2}$ (Elkies [10]), $3^{-1/2}$ (Bae [3], and Guy [16]), $\sqrt{3}\cdot(2\pi)^{-1/2}$ (Aliev [1]) and finally $\sqrt{2}\cdot\pi^{-1/2}$ (Dubroff, Fox, and Xu [9]).
A classic conjecture of Erdős states
Conjecture 1.3 (Erdős, \$ 500).
For a dissociated set $\mathcal{S}$, we have
$\left|\mathcal{S}(n)\right|\leq\log_{2}n+c$
where $c$ is a fixed constant.
This problem is listed as Problem #1\#1 in Bloom’s database of Erdős problems [4].
This note is mainly concerned with the greedy algorithm for generating these sets. It seems that this question hasn’t been discussed in literature before, although there are some interesting observations to be made. But before that, we will give a slightly generalized version of the theorem in [9] using a result of Harper [18]. The idea is to show that we can get better constants for a large number of $n$. To do so, we begin by defining
$B_{k}:=\binom{k}{\lfloor k/2\rfloor}\,=\,\left(\sqrt{\frac{2}{\pi}}+o(1)\right)\frac{2^{k}}{\sqrt{k}}$
and $U(n):=\max\{k:\ B_{k}\leq n\}$. The asymptote above is an application of Stirling’s approximation. For a sharper two sided bound, one can use the so called Robbins’ estimate [26]. Also, let $c_{\ast}:=\frac{1}{2}\log_{2}\left(\frac{\pi}{2}\right)$.
Theorem 1.4.
For any $c_{\ast}-1<C<c_{\ast}$, the set
$\mathcal{N}_{C}\,:=\,\left\{n\in\mathbb{N}:\,U(n)\,\leq\,\log_{2}n+\frac{1}{2}\log_{2}\log_{2}n+C\right\}$
has asymptotic density
$\mathbf{d}\left(\mathcal{N}_{C}\right)\,=\,2-2^{c_{\ast}-C}$
which is positive. In particular, for $C=0$, $\mathbf{d}\left(\mathcal{N}_{0}\right)=2-\sqrt{\frac{\pi}{2}}\approx 0.747$.
Proof.
Fix a large $k$ and consider some $n\in\mathcal{I}_{k}:=[B_{k},B_{k+1})$, and hence $U(n)=k$. Define
$\rho_{k}\,:=\,\frac{B_{k+1}}{B_{k}}$
and write $n=B_{k}r$ with $r\in[1,\rho_{k})$. Using $\rho_{k}=2(1+o(1))$, we may treat $r\in[1,2)$ asymptotically. Now, using Robbins’ estimate [26], we have
$\log_{2}n\,=\,\log_{2}B_{k}+\log_{2}r\,=\,k-\frac{1}{2}\log_{2}k-c_{*}+\log_{2}r+o(1)$
implying
$\log_{2}n+\frac{1}{2}\log_{2}\log_{2}n+C\,$ $=\,\left(k-\frac{1}{2}\log_{2}k-c_{*}+\log_{2}r\right)+\frac{1}{2}\log_{2}k+C+o(1)$
$=\,k+\log_{2}r+(C-c_{*})+o(1)$
since $\log_{2}\log_{2}n=\log_{2}k+o(1)$ uniformly for $n\in\mathcal{I}_{k}$.
So, the statement
$U(n)\,=\,k\,\leq\,\log_{2}n+\frac{1}{2}\log_{2}\log_{2}n+C$
is equivalent to $0\,\leq\,\log_{2}r+(C-c_{*})+o(1)\iff r\geq 2^{c_{*}-C}(1+o(1))$.
This selects the subinterval
$\left[B_{k}\cdot 2^{c_{*}-C}(1+o(1)),\,B_{k+1}\right)$
inside $\mathcal{I}_{k}$ whose relative size tends to
$\frac{\rho_{k}-2^{c_{*}-C}}{\rho_{k}-1}\longrightarrow 2-2^{c_{*}-C}$
which is also the limit of the asymptotic density since the lengths $|\mathcal{I}_{k}|$ grow geometrically.
2 The Greedy Algorithm
This section is concerned with the greedy algorithm of generating dissociated sets. It starts with two given integers $a>0$ and $b>a$. It sets $\gamma_{1}=a$, $\gamma_{2}=b$ and in the $r$-th step it chooses $\gamma_{r}$ to be the smallest integer greater than $\gamma_{r-1}$ such that the sequence $\{\gamma_{k}\}_{k=1}^{r}$ is dissociated.
Theorem 2.1.
For any given $a,b$, let $\Gamma:=\Gamma_{a,b}=\{\gamma_{1}=a,\gamma_{2}=b,\gamma_{3},\dots\}$ be the increasing dissociated sequence that the greedy algorithm produces. Then, there is an $n_{0}=n_{0}(a,b)$ such that
$\gamma_{n}=2\cdot\gamma_{n-1}$
for all $n\geq n_{0}$.
Before going to the proof, we need some definitions and a few preliminary lemmas. Define For a finite $A\subset\mathbb{N}$, we define
$\Sigma(A)\,:=\,\left\{\sum_{x\in T}x\,:\,T\subseteq A\right\}\subset\mathbb{N}\cup\{0\}$
to be the family of subset sums. We also define
$\Delta(A)\,:=\,\Sigma(A)-\Sigma(A)\,=\,\{u-v:\,u,v\in\Sigma(A)\}$
to be the difference set of subset sums. Note that if $d=u-v\in\Delta(A)$, then $d=v-u\in\Delta(A)$ and hence $\Delta(A)$ is symmetric. Also, for given integers $a$, $b>a$, we let $\Gamma_{a,b}=\{\gamma_{1}=a,\gamma_{2}=b,\gamma_{3},\dots\}$ denote the dissociated sequence that the greedy algorithm produces. We further denote $\Gamma_{a,b}(\kappa):=\Gamma_{a,b}\cap[1,\kappa]$ for some integer $\kappa$.
We also define
$B_{n}:=\left\{\sum_{i\in I}\gamma_{i}:\ I\subseteq[n]\right\}\,=\,\Sigma\Big(\Gamma(\gamma_{n})\Big)\subset\mathbb{Z}$
and
$D_{n}:=\left\{\sum_{i=1}^{n}\varepsilon_{i}\gamma_{i}:\ \varepsilon_{i}\in\{-1,0,1\}\right\}\,=\,\Delta\Big(\Gamma(\gamma_{n})\Big)\subset\mathbb{Z}$
to be the subset-sum set and the signed-sum set respectively. For a finite set $A$, we denote
$\mathscr{S}(A)\,:=\,\sum_{a\in A}a$
and $\mathscr{S}_{r}:=\mathscr{S}\Big(\Gamma(\gamma_{r})\Big)$. We also let $\mathscr{T}_{r}:=\mathscr{S}_{r}+1$.
Lemma 2.2.
Fix $n\geq 1$ and $x\in\mathbb{N}\setminus\Gamma(\gamma_{n})$. Then $\Gamma(\gamma_{n})\cup\{x\}$ is dissociated if and only if $x\notin D_{n}$.
Proof.
If $x\in D_{n}$, then
[…]
for some $\varepsilon_{i}\in\{-1,0,1\}$ and hence $\Gamma(\gamma_{n})\cup\{x\}$ is not dissociated.
Conversely, if $\Gamma(\gamma_{n})\cup\{x\}$ is not dissociated, then there exist
$\sum_{i=1}^{n}\varepsilon_{i}\gamma_{i}+\varepsilon_{x}x=0$
for some $\varepsilon_{i}\in\{-1,0,1\}$ for $i\leq n$ and $\varepsilon_{x}\in\{-1,0,1\}$, not all zero. If $\varepsilon_{x}=0$, this would be a forbidden nontrivial relation inside $\Gamma(\gamma_{n})$, contradicting that it is dissociated. Hence $\varepsilon_{x}=\pm 1$ and hence
$x\,=\,-\varepsilon_{x}\sum_{i=1}^{n}\varepsilon_{i}\gamma_{i}\,\in\,D_{n}$
hence completing the proof.
Lemma 2.3.
For every $n\geq 1$, $\mathscr{T}_{n}\notin D_{n}$. In other words, $\gamma_{n+1}\leq\mathscr{T}_{n}$ for all $n\geq 2$.
Proof.
This follows from the observation
$\left|\sum_{i=1}^{n}\varepsilon_{i}\gamma_{i}\right|\leq\sum_{i=1}^{n}|\varepsilon_{i}|\gamma_{i}\leq\mathscr{S}_{n}$
implying $D_{n}\subseteq[-\mathscr{S}_{n},\mathscr{S}_{n}]$.
Lemma 2.4.
For all $n\geq 1$, we have $\mathscr{T}_{n}\geq 2^{n}$.
Proof.
Follows from the fact that $\Gamma(\gamma_{n})$ is dissociated implying $|B_{n}|=2^{n}$ and $\Gamma$ increasing.
Equipped with these, we present the promised proof.
Proof of Theorem 2.1.
We proceed in three major steps. First, for $n\geq 1$, we define
$r_{n}\,:=\,\frac{\mathscr{T}_{n}}{2^{n}}\,\geq\,1$
by Lemma 2.4. Also, $\{r_{n}\}$ is non-increasing by Lemma 2.3.
Notice that there exists $N_{\ast}$ such that
$\gamma_{n+1}\,>\,\frac{1}{2}\mathscr{T}_{n}$
for all $n\geq N_{\ast}$. Indeed, if not, then for infinitely many $n$, we have
$\mathscr{T}_{n+1}=\mathscr{S}_{n}+\gamma_{n+1}+1\leq\mathscr{T}_{n}+\frac{1}{2}\mathscr{T}_{n}=\frac{3}{2}\mathscr{T}_{n}$
and hence
$r_{n+1}=\frac{\mathscr{T}_{n+1}}{2^{n+1}}\leq\frac{(3/2)\mathscr{T}_{n}}{2^{n+1}}=\frac{3}{4}r_{n}$
which implies $\{r_{n}\}$ goes to 0 over a subsequence, a contradiction.
The next step is to show that the greedy constraint forces large forbidden intervals. We will prove that $\{b+1,b+2,\dots,\gamma_{n}-1\}\subseteq D_{n-1}$ for all $n\geq 3$. We proceed by Induction. The base case is obvious since $\gamma_{3}$ is the smallest integer $b=\gamma_{2}$ not in $D_{2}$.
Now assume $\{b+1,\dots,\gamma_{n}-1\}\subseteq D_{n-1}$. Since $D_{n-1}\subseteq D_{n}$ (allowing an extra coordinate $\varepsilon_{n}=0$), we have $\{b+1,\dots,\gamma_{n}-1\}\subseteq D_{n}$. Also $\gamma_{n}\in D_{n}$ (take $\varepsilon_{n}=1$ and all other $\varepsilon_{i}=0$). Finally, by the greedy constraint, every integer $x$ with $\gamma_{n}<x<\gamma_{n+1}$ must lie in $D_{n}$. This completes the induction.
Finally, fix $n\geq 3$, $1\leq u\leq\gamma_{n}-b-1$ and $t:=\gamma_{n}-u$. Then, $b+1\leq t\leq\gamma_{n}-1$ and hence $t,-t\in D_{n-1}$ implying $u=\gamma_{n}+(-t)\in\gamma_{n}+D_{n-1}\subseteq D_{n}$ (using $\varepsilon_{n}=1$ plus a representation of t-t from the first n1n-1 terms).
This brings us to the third and final step where we prove $\gamma_{n+1}=2\gamma_{n}$ for all large $n$. Let’s choose $n$ large enough so that $D_{n-1}$ contains all integers $1,2,\dots,b$ and $2\gamma_{n}>\mathscr{S}_{n-1}$. Now,
$\big[\gamma_{n-1}-b-1\big]\subseteq D_{n-1}$
so that, if $\gamma_{n-1}\geq 2b+1$, then $\gamma_{n-1}-b-1\geq b$ implying $b]\subseteq D_{n-1}$. Since $\Gamma$ is strictly increasing and unbounded, there exists $N_{b}$ such that $\gamma_{n}\geq 2b+1$ for all $n\geq N_{b}$. Now, set $n_{0}:=\max\{N_{b}+1,\ N_{\ast}+1\}$. We claim that $\gamma_{n+1}=2\gamma_{n}$ for every $n\geq n_{0}$. So, from now on, fix an $n\geq n_{0}$.
Since $n-1\geq N_{b}$, we have $b]\subseteq D_{n-1}$ implying $\{b+1,\dots,\gamma_{n}-1\}\subseteq D_{n-1}$. Hence $\gamma_{n}-1]\subseteq D_{n-1}$. Therefore, for each $t\in[\gamma_{n}-1]$, we have $\gamma_{n}+t\in D_{n}$ (taking $\varepsilon_{n}=1$ and representing $t\in D_{n-1}$). So, $\big\{\gamma_{n}+1,\gamma_{n}+2,\dots,2\gamma_{n}-1\big\}\subseteq D_{n}$.
On the other hand, we will now show that $2\gamma_{n}\notin D_{n}$. If possible, let $2\gamma_{n}\in D_{n}$. Then
$2\gamma_{n}=\sum_{i=1}^{n}\varepsilon_{i}\gamma_{i}$
for $\varepsilon_{i}\in\{-1,0,1\}$. Let $\varepsilon:=\varepsilon_{n}\in\{-1,0,1\}$ and set
$m:=\sum_{i=1}^{n-1}\varepsilon_{i}\gamma_{i}\in D_{n-1}$
implying
$2\gamma_{n}=\varepsilon\gamma_{n}+m$
and hence $m=(2-\varepsilon)\gamma_{n}$.
If $\varepsilon=1$, then $m=\gamma_{n}$. But $\gamma_{n}\notin D_{n-1}$ because otherwise (by Lemma 2.2) adding $\gamma_{n}$ to $\Gamma(\gamma_{n-1})$ would not preserve the dissociated property, hence a contradiction.
If $\varepsilon=0$, then $m=2\gamma_{n}$. But any $m\in D_{n-1}$ satisfies $|m|\leq\mathscr{S}_{n-1}$ (using Lemma 2.3), while $2\gamma_{n}>\mathscr{S}_{n-1}$) since $n\geq n_{0}\geq N_{\star}+1$, a contradiction.
If $\varepsilon=-1$, then $m=3\gamma_{n}$ which is again impossible because $|m|\leq\mathscr{S}_{n-1}<2\gamma_{n}<3\gamma_{n}$.
This completes the proof.
The next natural question is to ask for an explicit $n_{0}(a,b)$. A clean upper bound follows essentially from the ideas discussed in the previous proof. So, we express this as a corollary.
Corollary 2.5.
We have
$n_{0}(a,b)\ \leq\ L+2K+5$
where $L:=\left\lceil\log_{2}(2b+2)\right\rceil$ and $K:=\left\lceil\log_{2}L\right\rceil$.
Proof.
Since $m\leq r=L+K+2$, we have
$n_{0}(a,b)\leq m+1+\left\lceil\log_{2}m\right\rceil$
and hence, it is enough to bound $\lceil\log_{2}m\rceil$ in terms of $K$. Since $m\leq r\leq L+K+2\leq 4L$ (using $K\leq L$ and $L\geq 2$), we have
$\log_{2}m\,\leq\,\log_{2}(4L)\,=\,2+\log_{2}L\leq 2+K$
and thus
$\left\lceil\log_{2}m\right\rceil\leq K+2$
so that substituting $m\leq L+K+2$ gives
$n_{0}(a,b)\leq(L+K+2)+1+(K+2)=L+2K+5$
as claimed.
3 Generalization I
The goal of this section and the next is to consider two different generalizations of the idea of a dissociated set.
Definition 3.1.
A set $\mathcal{S}=\{a_{1}<a_{2}<\dots\}\subset\mathbb{N}$ is said to be a $\mathcal{D}[g]$ set if
$r_{\mathcal{S}\cap[1,a_{m}]}(t)\,:=\,\left|\left\{I\subseteq[m]:\,\sum_{i\in I}a_{i}=t\right\}\right|\,\leq\,g$
for all $t\in\mathbb{Z}$ and $m\in\mathbb{N}$.
Under this terminology, a $\mathcal{D}[1]$ set is precisely the dissociated set that we have discussed before. We can upper bound $\mathcal{D}[g]$ sets using similar ideas as in [9].
Theorem 3.2.
Let $A={a_{1}<\cdots<a_{m}}\subset\mathbb{N}$ be a $\mathcal{D}[g]$ set. Then
$a_{m}\,\geq\,\frac{1}{g}\binom{m}{\lfloor\frac{m}{2}\rfloor}-1$
and hence $\binom{m}{\lfloor m/2\rfloor}\leq g(a_{m}+1)$.
Proof.
Define the Boolean cube
$V:=\left\{-\frac{1}{2},\frac{1}{2}\right\}^{m}$
and for $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{m})\in V$, define
$s(\varepsilon)\,:=\,\sum_{i\in I(\varepsilon)}a_{i}$
where $I(\varepsilon)\,:=\,\left\{i\in[m]:\varepsilon_{i}=\frac{1}{2}\right\}$. Now, take
$S:=\sum_{i=1}^{m}a_{i}$
and define the linear form
$L(\varepsilon)\,:$ $=\,\sum_{i=1}^{m}a_{i}\varepsilon_{i}$
$=\frac{1}{2}\sum_{i\in I(\varepsilon)}a_{i}-\frac{1}{2}\sum_{i\notin I(\varepsilon)}a_{i}$
$=\sum_{i\in I(\varepsilon)}a_{i}-\frac{1}{2}\sum_{i=1}^{m}a_{i}$
$=s(\varepsilon)-\frac{S}{2}$
and hence $L(\varepsilon)=L(\varepsilon^{\prime})\Longleftrightarrow s(\varepsilon)=s(\varepsilon^{\prime})$. So, each value of $L(\varepsilon)$ is attained by at most $g$ points $\varepsilon\in V$.
Now, define
$V_{-}:=\,\{\varepsilon\in V:\ L(\varepsilon)<0\}$
$V_{0}:=\,\{\varepsilon\in V:\ L(\varepsilon)=0\}$
$V_{+}:=\,\{\varepsilon\in V:\ L(\varepsilon)>0\}$
and note that $\varepsilon\mapsto-\varepsilon$ is a bijection $V_{-}\leftrightarrow V_{+}$ implying $|V_{-}|=|V_{+}|$. Choose any subset $V_{0}^{\prime}\subseteq V_{0}$ with
$|V_{0}^{\prime}|=2^{m-1}-|V_{-}|=\left\lceil\frac{|V_{0}|}{2}\right\rceil$
and define $F:=V_{-}\cup V_{0}^{\prime}$. Then, $|F|=2^{m-1}$.
Now define the boundary
[…]
so that $\eta$ is outside but within Hamming distance $1$ of $F$. We have
$|\partial F|\ \geq\ \binom{m}{\lfloor\frac{m}{2}\rfloor}$
using Harper’s vertex-isoperimetric inequality [18].
Fix $\eta\in\partial F$ and fix its corresponding $\xi\in F$. Therefore, we have
$L(\eta)-L(\xi)\,=\,\sum_{j=1}^{m}a_{j}(\eta_{j}-\xi_{j})\,=\,a_{i}(\eta_{i}-\xi_{i})\,\in\,\{+a_{i},-a_{i}\}$
since $\eta_{i}-\xi_{i}\in\{1,-1\}$. Now $L(\xi)\leq 0$ and $\eta\notin F$ implying $\eta\notin V_{-}\cup V_{0}^{\prime}$. So, $L(\eta)\geq 0$ with equality only if $\eta\in V_{0}\setminus V_{0}^{\prime}$. So every $\eta\in\partial F$ satisfies
$0\leq L(\eta)\leq a_{i}\leq a_{m}$
and all but those $\eta\in V_{0}\setminus V_{0}^{\prime}$ satisfy $0<L(\eta)\leq a_{m}$.
Also, we have
$\left|\partial^{+}F\right|:=\left\lvert\{\eta\in\partial F:\ L(\eta)>0\}\right\rvert\ \geq\ |\partial F|-g$
since $|V_{0}|=r_{A}(S/2)\cdot\mathbf{1}_{\{S\text{ is even}\}}\leq g$.
Now all values $L(\eta)$ for $\eta\in\partial^{+}F$ lie in the interval $(0,a_{m}]$, and any two distinct values differ by an integer because
$L(\eta)-L(\eta^{\prime})\,=\,L(\eta-\eta^{\prime})\,=\,\sum_{i=1}^{m}a_{i}(\eta_{i}-\eta^{\prime}_{i})$
and $\eta_{i}-\eta^{\prime}_{i}\in\{-1,0,1\}$, so the difference is an integer. Hence the set $\{L(\eta):\eta\in\partial^{+}F\}\subset(0,a_{m}]$ has at most $a_{m}$ distinct values (integer gaps $\geq 1$ inside an interval of length $a_{m}$). With multiplicity $\leq g$ per value, we have
$|\partial^{+}F|\leq g\cdot a_{m}$
implying
$a_{m}\geq\frac{|\partial F|}{g}-1\,\geq\,\frac{1}{g}\binom{m}{\lfloor\frac{m}{2}\rfloor}-1$
hence completing the proof.
Corollary 3.3.
Let $\mathcal{S}$ be a $\mathcal{D}[g]$ set. Then
$\left\lvert\mathcal{S}(n)\right\rvert\,\leq\,\log_{2}n\,+\,\frac{1}{2}\log_{2}\log_{2}n\,+\,\log_{2}g\,-\,\log_{2}\left(\sqrt{\frac{2}{\pi}}\right)\,+\,o(1)$
where $\mathcal{S}(n):=\mathcal{S}\cap[1,n]$.
Remark: It should be briefly noted that there are some cute consequences of Corollary 3.3. For instance, it immediately implies that for a given $k$ and $\ell$, there are infinitely many $N$ that can be written as the sum of distinct $k$-th powers in at least $\ell$ different ways. Similarly, there are infinitely many $N\in\mathbb{N}$ that can be written as the sum of primes (on your favourite arithmetic progression) in $\ell$ different ways.
On the other hand, the greedy algorithm for $\mathcal{D}[g]$ sets also exhibit the same doubling property as discussed, with the proof being very similar to that of Theorem 2.1. So, instead of giving the full proof again, we will only provide an outline.
Theorem 3.4.
For any given $a,b$, let $\gamma_{1}=a$, $\gamma_{2}=b$ and in let $\gamma_{r}$ to be the smallest integer greater than $\gamma_{r-1}$ such that the sequence $\{\gamma_{k}\}_{k=1}^{r}$ is $\mathcal{D}[g]$. Let $\Gamma:=\Gamma_{a,b}^{(g)}=\{\gamma_{1}=a,\gamma_{2}=b,\gamma_{3},\dots\}$. Then, there is an $n_{0}=n_{0}(a,b,g)$ such that
$\gamma_{n}=2\cdot\gamma_{n-1}$
for all $n\geq n_{0}$.
Proof.
As advertised, the proof is formally identical to the dissociated ($g=1$) proof once one replaces the signed–sum set $D_{n}$ by the forbidden–shift condition expressed via the representation function
$r_{n}(s)\,:=\,\left|\left\{I\subseteq[n]\,:\,\sum_{i\in I}\gamma_{i}=s,\right\}\right|\,=\,\big[z^{s}\big]\prod_{i=1}^{n}\big(1+z^{\gamma_{i}}\big)$
for $s\in\mathbb{Z}$.
Adding $x$ transforms $r_{n}$ by the standard subset-sum recurrence
$r_{n+1}(s)=r_{n}(s)+r_{n}(s-x)$
and hence $\Gamma\big(\gamma_{n}\big)\cup\{x\}$ is $\mathcal{D}[g]$ iff $r_{n}(s)+r_{n}(s-x)\leq g$ for all $s$ iff $x$ is not in the forbidden set
$F_{n}\,:=\,\left\{x\in\mathbb{N}:\exists s\ \text{with }r_{n}(s)+r_{n}(s-x)\geq g+1\right\}$
which will act as the analogue of $D_{n}$ from the previous proof.
Finally, borrowing notations from the previous proof, we have
$2^{n}\,=\,\sum_{s}r_{n}(s)\,\leq\,g(\mathscr{S}_{n}+1)\,=\,g\mathscr{T}_{n}$
so the potential becomes
$r_{n}^{(g)}\,:=\,\frac{g\mathscr{T}_{n}}{2^{n}}$
and the rest of the proof runs through.
4 Generalization II
Definition 4.1.
A set $\mathcal{S}\subset\mathbb{N}$ is called a $\mathcal{D}_{k}$ set if any equality of the form
$\sum_{s\in\mathcal{S}}\varepsilon_{s}\cdot s=0$
where $\varepsilon_{s}\in\{-k,\dots,-1,0,+1,\dots,k\}$ implies that all the $\varepsilon_{s}$’s are 0.
Dissociated sets in this language are $\mathcal{D}_{1}$ sets and they have been called detecting sets in literature [21, 22]. In this case again, we can get an upper bound by generalizing the argument in [9]. This is what we will now exhibit.
Theorem 4.2.
Let $\mathcal{S}=\{a_{1}<\cdots<a_{m}\}\subset\mathbb{N}$ be a $\mathcal{D}_{k}$ set. Then, we have
$a_{m}\,\geq\,(1+o(1))\,\frac{(k+1)^{m}}{2}\,\sqrt{\frac{6}{m\pi k\,(k+2)}}$
and hence, for an infinite $\mathcal{D}_{k}$ set $\mathcal{S}$, we have
$\left\lvert\mathcal{S}(n)\right\rvert\,\leq\,\log_{k+1}n\,+\,\frac{1}{2}\log_{k+1}\log_{2}n\,+\,\frac{1}{2}\log_{k+1}\left(\frac{2\pi}{3}\,k(k+2)\right)\,+\,o(1)$
where $\mathcal{S}(n)\,:=\,\mathcal{S}\cap[1,n]$.
Proof.
As before, we again define
$V:=[0,k]^{m}=\{0,1,\dots,k\}^{m}$
viewed as the vertex set of the grid graph
$G=\mathop{{{{\times}}}}_{i\in[m]}P_{k+1}$
expressed as the Cartesian product of path graphs - two vectors are adjacent iff they differ by $\pm 1$ in exactly one coordinate. For $x=(x_{1},\dots,x_{m})\in V$, define
$s(x):=\sum_{i=1}^{m}a_{i}x_{i}\in\mathbb{Z},\quad S:=\sum_{i=1}^{m}a_{i},\quad L(x):=s(x)-\frac{k}{2}S$
and note $L(k\cdot\mathbf{1}-x)=-L(x)$ where $\mathbf{1}=(1,\dots,1)$. Let
$c_{m,k}(t)\,:=\,\big[x^{t}\big]\,\Big(1+x+\cdots+x^{k}\Big)^{m}$
be the number of vectors $x\in V$ with $|x|:=x_{1}+\cdots+x_{m}=t$. The sequence $t\mapsto c_{m,k}(t)$ is symmetric and unimodal, and it attains its maximum at $t=\lfloor\frac{km}{2}\rfloor$.
Now, note that $\mathcal{S}$ is a $\mathcal{D}_{k}$ set, then
$s(x)=s(y)\implies 0=s(x)-s(y)=\sum_{i=1}^{m}(x_{i}-y_{i})a_{i}$
with $x_{i}-y_{i}\in\{-k,\dots,k\}$ and hence, the map $s:V\to\mathbb{Z}$, $x\mapsto\sum a_{i}x_{i}$ is injective. In particular, $L$ is also injective.
Again as before, we set
$V_{-}\,:=\,\{x\in V:L(x)<0\}$
$V_{0}\,:=\,\{x\in V:L(x)=0\}$
$V_{+}\,:=\,\{x\in V:L(x)>0\}$
and use $L(k\mathbf{\cdot}1-x)=-L(x)$ to get $|V_{-}|=|V_{+}|$.
Again, choose $V_{0}^{\prime}\subseteq V_{0}$ so that
$\big|F\big|\,:=\,\big|V_{-}\big|\,+\,\big|V_{0}^{\prime}\big|\,=\,\left\lfloor\frac{|V|}{2}\right\rfloor$
and hence, $F\subseteq V$ has size $\left\lfloor\frac{(k+1)^{m}}{2}\right\rfloor$. Finally, also define the vertex boundary $\partial F$ to be the set of all $y\in V\setminus F$ such that there is $x\in F$ adjacent to $y$ in the grid.
From Theorem 11 in [19] (also see [31]), we have
$|N(F)\setminus F|\,=\,|N(F)|-|F|\,\geq\,|N(C)|-|C|$
implying $|\partial F|\,\geq\,|\partial C|$.
On the other hand, let $t_{0}:=\left\lfloor\frac{km}{2}\right\rfloor$ and $L_{t}:=\{x\in V:|x|=t\}$) so that $|L_{t}|=c_{m,k}(t)$. Because the level sizes are symmetric and unimodal with maximum at $t_{0}$ (see [20] for example), we have
$\sum_{t<t_{0}}\big|L_{t}\big|\,\leq\,\frac{\big|V\big|-\big|L_{t_{0}}\big|}{2}$
so that when you take any initial segment $C$ of size $\left\lfloor\frac{|V|}{2}\right\rfloor$, you must include at most half of the middle layer $L_{t_{0}}$, i.e., $|C\cap L_{t_{0}}|\leq\frac{|L_{t_{0}}|}{2}$.
But, every vertex in $L_{t_{0}}\setminus C$ has a neighbor in $L_{t_{0}-1}\subseteq C$ and hence
$L_{t_{0}}\setminus C\,\subseteq\,\partial C$
implying
$\big|\partial F\big|\,\geq\,\big|\partial C\big|\,\geq\,\big|L_{t_{0}}\setminus C\big|\,\geq\,\big|L_{t_{0}}\big|-\frac{|L_{t_{0}}|}{2}\,=\,\frac{1}{2}\,\big|L_{t_{0}}\big|\,=\,\frac{1}{2}\,M_{m,k}$
where $M_{m,k}$ is the maximum of $c_{m,k}(t)$ over $t$.
Now, consider any $y\in\partial F$ with $L(y)>0$. By definition of boundary, there exists an adjacent $x\in F$ differing from $y$ in exactly one coordinate $i$ by $\pm 1$. Since $x\in F$ implies $L(x)\leq 0$ and $L(y)>0$, necessarily $L(y)-L(x)=+a_{i}$. Hence
$0\,<\,L(y)\,=\,L(x)+a_{i}\,\leq\,a_{i}\,\leq\,a_{m}$
and hence every such $y$ yields a value $L(y)\in(0,a_{m}]$.
Moreover, for any $y,y^{\prime}\in V$, we have
$L(y)-L(y^{\prime})=s(y)-s(y^{\prime})\in\mathbb{Z}$
and hence all $L$-values lie in a single coset of $\mathbb{Z}$ and are spaced by integers. Since $L$ is injective, distinct $y$’s give distinct $L(y)$’s. Therefore the set
$\big\{L(y):y\in\partial F,\ L(y)>0\big\}\subset(0,a_{m}]$
has size at most $a_{m}$. Finally, because $|V_{0}|\leq 1$, at most one boundary vertex can have $L(y)=0$. Hence
$a_{m}\,\geq\,\big|\partial F\big|-1\,\geq\,\frac{1}{2}\,M_{m,k}-1$
since $|\partial F|\,\leq\,a_{m}+1$.
By unimodality of $M_{m,k}$ and Central Limit Theorem, we have (see [20]and [9])
$M_{k,m}\sim(k+1)^{m}\sqrt{\frac{6}{m\pi k\,(k+2)}}$
and equivalently, the $(k+1)$-multinomial maximum $\sup_{t}\binom{m,k+1}{t}$ satisfies the same estimate. This completes the proof.
Of course one could give similar generalizations of Theorem 4.2 or Corollary 3.3 along the lines of Theorem 1.4. Also, the question of finding $n_{0}(a,b,g)$ from 3.4 can be handled similarly as done after Theorem 2.1. But the real question right now is understanding the greedy algorithm for $\mathcal{D}_{k}$ sets. Unfortunately, this seems to be a much trickier problem than the two similar ones we have handled before. From numerical evidence, it can be easily conjectured that
$\gamma_{r+1}^{(k)}\,=\,(k+1)\cdot\gamma_{r}^{(k)}$
for all $n>n_{0}(a,b,k)$. Using similar techniques as in the proof of Theorem 2.1, it is not too difficult to show
$\lim_{n\to\infty}\frac{\gamma_{n+1}^{(k)}}{\gamma_{n}^{(k)}}\to(k+1)$
but going the next step looks very tricky even for specific small values of $k$. The resolution of this will be left to the interested readers.
5 Dissociated Sets Avoiding Geometric Progression
Finally, we will conclude by making an observation. From Theorems 2.1 and 3.3, it might look like dissociated sets must contain lots of geometric progressions. We will show that this is probably not true by proving it in a large collection of dissociated sets.
To begin with, let $\mathscr{D}(n)$ be the collection of all dissociated subsets of $[n]$. From Theorem 1.2, for $M:=\lfloor 2\log_{2}n\rfloor$, we have
$\mathscr{D}(n)\,\leq\,\sum_{j=0}^{M}\binom{n}{j}\,\ll\,(M+1)\binom{n}{M}\,\ll\,(M+1)\left(\frac{en}{M}\right)^{M}$
implying $\mathscr{D}(n)\leq\exp\Big(\mathcal{O}\big((\ln n)^{2}\big)\Big)$.
Now, fix a large $n$ and consider the following family. Set $\tilde{m}:=\lfloor\log_{2}n\rfloor$ and $m:=\tilde{m}-L$ for an integer $L=\mathcal{O}(\log_{2}\log_{2}n)$. For each $t\in\{0,1,\dots,m-1\}$, define the dyadic interval
$I_{t}:=\left(\frac{n}{2^{t+1}},\,\frac{n}{2^{t}}\right]\cap\big(2\mathbb{Z}+1\big)$
and define $\mathcal{F}_{L}(n)$ to be the family of sets
$S(\mathbf{q})\,=\,\big\{2^{t}q_{t}:0\leq t\leq m-1\big\}\subset[n]$
where $q_{t}\in I_{t}$ for each $t$. Now, let
$\mathscr{G}_{\mathcal{F}}(n):=\#\big\{S\in\mathcal{F}_{L}(n):\nexists\,a,b,c\in S:b^{2}=ac\big\}$
and $\mathscr{D}_{\mathcal{F}}(n):=|\mathcal{F}_{L}(n)|$.
Now, every $\mathcal{S}\in\mathcal{F}_{L}(n)$ is dissociated and $|\mathcal{S}|\sim(\log_{2}n)$. Indeed, if
$\sum_{t=0}^{m-1}\varepsilon_{t}\big(2^{t}q_{t}\big)=0$
with $\varepsilon_{t}\in\{-1,0,1\}$, then
$\varepsilon_{t_{0}}\,2^{t_{0}}\,q_{t_{0}}\,\equiv\,0\pmod{2^{t_{0}+1}}$
where $t_{0}$ be the smallest index with $\varepsilon_{t_{0}}\neq 0$. But $q_{t_{0}}$ is odd, so
$\varepsilon_{t_{0}}2^{t_{0}}q_{t_{0}}\equiv\pm 2^{t_{0}}\not\equiv 0\pmod{2^{t_{0}+1}}$
which is a contradiction, proving $\varepsilon_{t}=0$ for all $t$.
On the other hand, we also have
$\log_{2}\big|\mathcal{F}_{L}(n)\big|\,=\,\sum_{t=0}^{m-1}\log_{2}\big|I_{t}\big|\,=\,\sum_{t=0}^{m-1}\Big(\log_{2}n-t+O(1)\Big)\,=\,\frac{m^{2}}{2}+\mathcal{O}\big(m\log_{2}m\big)$
implying
$\big|\mathcal{F}_{L}(n)\big|\,=\,2^{\frac{1}{2}(\log_{2}n)^{2}+o\Big((\log n)^{2}\Big)}$
and hence, the family $\mathcal{F}_{L}(n)$ is quite large. Notice that this also proves $\mathscr{D}(n)=\exp\Big(\Theta\big((\ln n)^{2}\big)\Big)$.
Now, write $x_{t}:=2^{t}q_{t}$ and notice that if $x_{j}^{2}=x_{i}x_{k}$, then $2j=i+k$ and $q_{j}^{2}=q_{i}q_{k}$. This implies that if we set $i,j,k$ with $i+k=2j$, then
$\mathbb{P}\big(x_{j}^{2}=x_{i}x_{k}\big)\,\leq\,\frac{1}{|I_{k}|}$
since at most one choice of $q_{k}\in I_{k}$ can satisfy $q_{k}=q_{j}^{2}/q_{i}$.
Now, for a fixed $k$, we have $\big\{(i,j):i<j<k,\ i+k=2j\big\}\,\leq\,k/2$. Therefore, for $k\leq m-1$, we have $|I_{k}|\gg n\,2^{-k}$. So, if $\mathcal{S}$ is chosen from $\mathcal{F}_{L}(n)$ uniformly randomly, then
$\mathbb{P}\big(\exists\,a,b,c\in\mathcal{S}:b^{2}=ac\big)\,\ll\,\sum_{k=2}^{m-1}\left(\frac{k}{2}\right)\frac{2^{k}}{n}\ll\tilde{m}\,2^{-L}$
and consequently
$\frac{\mathscr{G}_{\mathcal{F}}(n)}{\mathscr{D}_{\mathcal{F}}(n)}\,\geq\,1-\mathcal{O}\Big(\tilde{m}\,2^{-L}\Big)$
thus showing that most members of $\mathcal{F}_{L}(n)$ don’t contain a three term geometric progression.
Remark: We can sharpen the result to
$\frac{\mathscr{G}_{\mathcal{F}}(n)}{\mathscr{D}_{\mathcal{F}}(n)}\,\geq\,1-\mathcal{O}\left(\frac{1}{\big(\log_{2}n\big)^{A+1}}\right)$
by choosing $L:=\left\lceil(A+2)\log_{2}m\right\rceil$ and doing a similar calculation.
6 Acknowledgements
I would like to thank Prof. Ramachandran Balasubramanian, Prof. Greg Martin, Prof. Ilya Shkredov, Prof. Sándor Kiss, Satvik Saha and Sohom Gupta for the discussions I had with them about this topic.